I am stuck! Please help :)
Solve the formula for the specified variable.
1. P=A/1+rt Solve for r

Answer by mananth(16946)

Solve the formula for the specified variable.
1. P=A/1+rt Solve for r

1+rt= A/P
rt= A/P -1
r= A/P-1 /t